OpenStudy (anonymous):

A sample of C3H8 has 3.28 × 1024 H atoms. How many carbon atoms does the sample contain? What is the total mass of the sample in grams?

OpenStudy (anonymous):

It should be 3.28 X 10^24 From this find the moles of Hydrogen by applying this formula. moles = no. of atoms/ Avogadro's no (NA) moles = 5.446 then use the ratio of Carbon and Hydrogen from the formula: C = H 3 = 8 3/8 * 5.466 = 5.466 2.05 moles of C. now: no of C atoms = moles X NA = 2.05 X 6.022 X 10^23 = 1.27 X10^24 2nd Part: One mole of C3H8 = 3 moles of C 1/3 * 2.05 = 2.05 moles of C 0.68 moles of C3H8. mass of C3H8 = moles X molar mass mass of C3H8 = 0.68 X 44 = 30.2 grams of C3H8